### v2

# Orbital orders under magnetic fields in cubic PrIr2Zn20

Abstract
We study orbital orders in quadrupolar heavy-fermion compound PrIr2Zn20 under magnetic fields on the basis of the Landau theory. Assuming Eg orbital (electric quadrupolar) orders in the cubic symmetry with the ordering wavenumber at the L points in the face-centered cubic lattice Brillouin zone as observed experimentally, we construct the Landau free energy and analyze it. We find that the unidentified high-temperature ordered phase under the magnetic field $\bm{H}\parallel[001]$ reported earlier can be interpreted as the consequence of the internal rotation in the orbital moments of f electrons at the Pr site. We discuss the phase diagram for various magnetic-field directions and also possible double-$\bm{q}$ quadrupolar orders in this system.
pacs:
Valid PACS appear here
I Introduction
Strongly correlated electron systems with orbital moments have attracted significant attention in recent years. Tokura and Nagaosa (2000) Various multipolar orders appear in d- and f-electron-based materials and have been intensively studiedMydosh and Oppeneer (2014); Kuramoto et al. (2009); Chen et al. (2010). Among them, electric quadrupole orders are the most typical and extensively investigatedShiina et al. (1997); Onimaru et al. (2005); Chubukov et al. (2016). In addition, spin nematic orders have been also studied as well.Andreev and Grishchuk (1984); Läuchli et al. (2006); Tsunetsugu and Arikawa (2006); Jiang et al. (2023) The nature of orbital degrees of freedom leads to spatially anisotropic interactions, making the symmetry in orbital space generally anisotropic. Each type of orbital order has unique properties, and understanding these properties remains a key issue in condensed matter theory and experiment.
[…]
Among the 1-2-20 compounds, PrIr2Zn20 is a prototypical material, exhibiting quadrupole ordering below 1 K and undergoing a phase transition to a superconducting state at 0.1 KOnimaru et al. (2011); Ishii et al. (2011); Onimaru et al. (2016). The early neutron scattering experimentIwasa et al. (2017) under an applied magnetic field $\bm{H}\parallel[\bar{1}10]$ suggested an ordering vector of $\bm{q}=(\pi,\pi,\pi)$, corresponding to the L point at the first Brillouin zone boundary. Here, we have set the lattice constant to unity. The analysis of induced magnetic moments indicates an orbital order of the $x^{2}-y^{2}$ type for $\bm{H}\parallel[\bar{1}10]$. Recent reinvestigations of the temperature–magnetic field phase diagram reveal strong anisotropy and highlight the existence of the high-temperature phaseKittaka et al. (2024) reported in earlier studiesOnimaru et al. (2011) across a wide range of magnetic field directions from [001] to [111], while it is absent for $\bm{H}\parallel[110]$.
In this paper, we analyze a phenomenological Landau model for PrIr2Zn20, incorporating $E_{g}$ quadrupole moments at the L points. We elucidate the nature of $E_{g}$ orbital ordering under magnetic fields and its dependence on field direction. We examine the order parameter suggested by the neutron scattering experiments for $\bm{H}\parallel[\bar{1}10]$, construct phase diagrams under applied magnetic fields, and compare them with experimental results. The ordering vectors at the L points reveal unique multiple-$\bm{q}$ orders, which differs from the previous studies about quadrupole orders at the X point ordering vectors in the face-center cubic lattice.Tsunetsugu et al. (2021); Hattori et al. (2023)
This paper is organized as follows. In Sec. II, we construct the Landau free energy for the $E_{g}$ quadrupole moments at wave vectors $\bm{q}=(\pi,\pi,\pi)$ and $(0,0,0)$. In Sec. III, we present the results of the Landau analysis for single-$q$ orbital orders, outlining the fundamental properties of the system. We then discuss multiple-$q$ orbital orders and the stability of various ordered phases in Sec. IV. In Sec. V, we compare our results with experimental data for PrIr2Zn20 and related materials. Finally, Sec. VI provides a summary of this paper.
II Model
We start to introduce the exchange interactions between the Eg orbital moments in Sec. II.1 and the Landau free energy for the order parameters at the L points in Sec. II.2. Mode coupling terms related to the order parameter at the L points and other degrees of freedom are discussed in Sec. II.5
Figure 1: Fourier transform of the exchange interaction for $J_{1}=1$ and (a) $(J_{2},J_{3},J_{4})=(0.0,0.0,0.0)$, (b) $(0.0,0.0,1.0)$, and (c) $(0.5,0.3,0.4)$. (d) The first Brillouin zone of the fcc lattice. The four L points are indicated by filled red circles and other symmetric points are indicated by filled black circles.
II.1 Exchange interaction
Let us denote Eg quadrupole moments at the position $\bm{r}=(x,y,z)$ in the diamond structure as $\bm{\phi}_{\bm{r}}^{\alpha}=(u_{\bm{r}}^{\alpha},v_{\bm{r}}^{\alpha})$, where $\alpha=$A or B is the sublattice index. We take one of the A sublattice site at the origin $\bm{r}=(0,0,0)$ and its pair B sublattice site in the unit cell at $\bm{r}+(1,1,1)/4$. $\bm{\phi}_{\bm{r}}^{\alpha}$ form a two-dimensional vector field as a classical variable. Here, $(u,v)$ represents $(u,v)\sim\{2z^{2}-x^{2}-y^{2},\sqrt{3}(x^{2}-y^{2})\}$, where $\bm{r}=(x,y,z)$ is the real space coordinate. The Fourier transform in the wavenumber ${\bm{k}}$ space is defined as
${\boldsymbol{\phi}}_{{\bm{k}}}^{\alpha}=\begin{bmatrix}u^{\alpha}_{{\bm{k}}}\\ v^{\alpha}_{{\bm{k}}}\end{bmatrix}=\frac{1}{\sqrt{N}}\sum_{\bm{r}}e^{-i{\bm{k}}\cdot\bm{r}}\begin{bmatrix}u^{\alpha}_{\bm{r}}\\ v^{\alpha}_{\bm{r}}\end{bmatrix}.$ (1)
Here, the summation for ${\bm{r}}$ runs over the A-sublattice sites in the diamond structure and $N$ is the number of the unit cell.
The exchange Hamiltonian in the diamond structure is in general written as
$H=\sum_{{\bm{k}}}[u^{A}_{-{\bm{k}}},v^{A}_{-{\bm{k}}},u^{B}_{-{\bm{k}}},v^{B}_{-{\bm{k}}}]\begin{bmatrix}H^{AA}_{{\bm{k}}}&H^{AB}_{{\bm{k}}}\\ {H^{AB}_{{\bm{k}}}}^{\dagger}&H^{AA}_{{\bm{k}}}\end{bmatrix}\begin{bmatrix}u^{A}_{{\bm{k}}}\\ v^{A}_{{\bm{k}}}\\ u^{B}_{{\bm{k}}}\\ v^{B}_{{\bm{k}}}\end{bmatrix},$ (2)
where $H^{\alpha\beta}_{{\bm{k}}}$ is a $2\times 2$ matrix. We focus on the orders with the ordering wavenumber at the L points: ${\bm{k}}_{1}\equiv(\pi,\pi,\pi)$, ${\bm{k}}_{2}\equiv(-\pi,\pi,\pi)$, ${\bm{k}}_{3}\equiv(-\pi,-\pi,\pi)$, and ${\bm{k}}_{4}\equiv(\pi,-\pi,\pi)$. Since the exchange interactions at the L points are isotropic due to the symmetry, we retain such exchange interactions while neglect anisotropic ones. Although the latter exchange interactions are generally present and can influence the ordering vectors and the excitation energy of quadrupole wave, it is no effect on the transition temperature for the order at the L points. Thus, a model without anisotropic interactions can be a good approximation as a minimal model. In this section, we consider exchange interactions up to the fourth neighbor interactions, while we will simplify the model in the later sections. Hereafter, the lattice constant $a$ is set to $a=1$.
The next nearest-neighbor and the fourth-neighbor bonds are those between the A-A (B-B) sublattices. They are
$H^{AA}_{{\bm{k}}}$ $=[J_{2}\gamma_{2}({\bm{k}})+J_{4}\gamma_{4}({\bm{k}})]\hat{\sigma}_{0}$ (3)
Here, $\hat{\sigma}_{0}={\rm diag}(1,1)$ and $\hat{\sigma}^{x,z}$ is the Pauli matrix and we have followed the phase convention in Ref. [Hattori2014]. $J_{i}$’s are isotropic coupling constants. The expressions for $\gamma$’s are listed in Table 1.
The nearest-neighbor and the third-neighbor bonds are those between the A-B sublattices. For the nearest-neighbor interaction, there is no anisotropic term owing to the symmetry. The form is thus given as
$H^{AB}_{{\bm{k}}}$ $=\frac{1}{2}[J_{1}\gamma_{1}({\bm{k}})+J_{3}\gamma_{3}({\bm{k}})]\hat{\sigma}_{0}.$ (4)
We note that $\gamma_{1}$ and $\gamma_{3}$ are complex. From the neutron experimental data, the ordering patterns are consistent with those expected for $J_{1}>0$.
For describing the symmetry breaking at the L point modulation within our model, it is important to check which interactions are important. As will be explained later, four special wavenumbers at the $\Gamma$ point and the X points $\{(2\pi,0,0),(0,2\pi,0),(0,0,2\pi)\}$ are important other than the L point. First, among the same sub-lattice interactions, we note that the next-nearest interaction $\gamma_{2}$ vanishes at the four L points, while the fourth neighbor one $\gamma_{4}$ is finite as shown in Table 1. This means that positive $J_{4}$ favors the order at the L points while does not those at the $\Gamma$ and X points. For the inter sublattice ones, the nearest-neighbor interaction $\gamma_{1}$ and the third-neighbor one $\gamma_{3}$ are finite, while both vanish at the X points. In this sense, the effects of $J_{1}$ and $J_{3}$ are similar. Diagonalizing the matrix in Eq. (2), one obtains the eigenvalues $J({\bm{k}})$. Figure 1 shows $J({\bm{k}})$ for typical parameter sets with $J_{1}=1$ as a unit of energy. For the nearest-neighbor model with $J_{4}=0$ in Fig. 1(a), $J({\bm{k}})$ is smallest at the $\Gamma$ point, leading to a Néel type antifero quadrupole order. When $J_{4}$ is sufficiently large or $J_{2}$ is positive and not too large, the $J({\bm{k}})$ has minima at the L points as shown in Figs. 1(b) and 1(c), and these cases are situations we analyze in this paper.
Table 1: exchange interaction coefficients. The values of $\gamma({\bm{k}})$’s at ${\bm{k}}=(0,0,0),(0,0,2\pi)$, and $(\pi,\pi,\pi)$ are also shown. $\eta_{{\bm{k}}}\equiv e^{i(k_{x}+k_{y}+k_{z})/4}$, $\zeta^{\pm}_{{\bm{k}}}\equiv\pm 2(\cos\frac{k_{x}}{2}+\cos\frac{k_{y}}{2}+\cos\frac{k_{z}}{2})-3$. The mean-field second-order transition temperatures $T_{c}$’s are $T_{c}=2|J_{1}-3J_{3}|+6J_{4}$ for the L points, $T_{c}=4|J_{1}+3J_{3}|-12J_{2}-6J_{4}$ for the $\Gamma$ points, and $T_{c}=4J_{2}-6J_{4}$ for the X points.
notation expression $\Gamma(0,0,0)$ X(0,0,$2\pi$) L($\pi,\pi,\pi$)
$\gamma_{2}({\bm{k}})$ $2\left(\cos\frac{k_{x}}{2}\cos\frac{k_{y}}{2}+\cos\frac{k_{y}}{2}\cos\frac{k_{z}}{2}+\cos\frac{k_{z}}{2}\cos\frac{k_{x}}{2}\right)$ $6$ $-2$ 00
$\gamma_{4}({\bm{k}})$ $\cos k_{x}+\cos k_{y}+\cos k_{z}$ $3$ $3$ $-3$
$\gamma_{1}({\bm{k}})$ $4\eta_{{\bm{k}}}\left(\cos\frac{k_{x}}{4}\cos\frac{k_{y}}{4}\cos\frac{k_{z}}{4}+i\sin\frac{k_{x}}{4}\sin\frac{k_{y}}{4}\sin\frac{k_{z}}{4}\right)$ $4$ 00 $-2$
$\gamma_{3}({\bm{k}})$ $4\eta_{{\bm{k}}}\left(\cos\frac{k_{x}}{4}\cos\frac{k_{y}}{4}\cos\frac{k_{z}}{4}\zeta^{+}_{{\bm{k}}}+i\sin\frac{k_{x}}{4}\sin\frac{k_{y}}{4}\sin\frac{k_{z}}{4}\zeta^{-}_{{\bm{k}}}\right)$ $12$ 00 $6$
II.2 Landau Expansion
We now focus on the order parameter at the four equivalent L points. Note that ${\bm{k}}_{1}+{\bm{k}}_{2}+{\bm{k}}_{3}+{\bm{k}}_{4}\equiv{\bf 0}$ and that the fields at the L points are real since ${\bm{k}}_{n}\equiv-{\bm{k}}_{n}$.
The Landau’s free energy $F$ for the quadrupole moments is written as
$F$ $=\sum_{{\bm{k}}}[{\boldsymbol{\phi}}^{A}_{-{\bm{k}}},{\boldsymbol{\phi}}^{B}_{-{\bm{k}}}]\left[\frac{T}{2}+\hat{J}({\bm{k}})\right]\begin{bmatrix}{\boldsymbol{\phi}}^{A}_{{\bm{k}}}\\ {\boldsymbol{\phi}}^{B}_{{\bm{k}}}\end{bmatrix}$
$+\sum_{\alpha}\Big{[}-b\sum_{{\bm{k}},{\bm{k}}^{\prime},{\bm{k}}^{\prime\prime}}u^{\alpha}_{{\bm{k}}}\Big{(}u^{\alpha}_{{\bm{k}}^{\prime}}u^{\alpha}_{{\bm{k}}^{\prime\prime}}-3v^{\alpha}_{{\bm{k}}^{\prime}}v^{\alpha}_{{\bm{k}}^{\prime\prime}}\Big{)}\delta_{{\bm{k}}+{\bm{k}}^{\prime}+{\bm{k}}^{\prime\prime},{\bm{G}}}$
$+\frac{c}{4}\sum_{{\bm{k}},{\bm{k}}^{\prime},{\bm{k}}^{\prime\prime},{\bm{k}}^{\prime\prime\prime}}({\boldsymbol{\phi}}^{\alpha}_{{\bm{k}}}\cdot{\boldsymbol{\phi}}^{\alpha}_{{\bm{k}}^{\prime}})({\boldsymbol{\phi}}^{\alpha}_{{\bm{k}}^{\prime\prime}}\cdot{\boldsymbol{\phi}}^{\alpha}_{{\bm{k}}^{\prime\prime\prime}})\delta_{{\bm{k}}+{\bm{k}}^{\prime}+{\bm{k}}^{\prime\prime}+{\bm{k}}^{\prime\prime\prime},{\bm{G}}}\Big{]}.$ (5)
Here, $\hat{J}({\bm{k}})$ is the $4\times 4$ matrix in Eq. (2) with its eigenvalue $J({\bm{k}})$ and $T$ is the temperature and $\bm{G}$ is a reciprocal lattice vector. We have taken into account local cubic and fourth-order terms only, which arise from local anisotropy such as CEF potential.
II.3 L point free energy
When squeeze out the L point components ${\boldsymbol{\phi}}^{\alpha}_{n}=(u_{n}^{\alpha},v_{n}^{\alpha})\equiv{\boldsymbol{\phi}}^{\alpha}_{{\bm{k}}_{n}}(n=1,2,3,4)$ with the short handed notation, one finds
$F_{\rm L}$ $=\frac{a_{\rm L}}{2}\sum_{\alpha,n}|{\boldsymbol{\phi}}_{n}^{\alpha}|^{2}+J_{\rm L}\sum_{n}{\boldsymbol{\phi}}^{A}_{n}\cdot{\boldsymbol{\phi}}^{B}_{n}+\frac{c}{4}\sum_{\alpha,n}|{\boldsymbol{\phi}}_{n}^{\alpha}|^{4}$
$+\frac{c^{\prime}}{4}\sum_{\alpha}\sum_{m<n}|{\boldsymbol{\phi}}^{\alpha}_{m}|^{2}|{\boldsymbol{\phi}}^{\alpha}_{n}|^{2}+\frac{c^{\prime\prime}}{4}\sum_{\alpha}\sum_{m<n}({\boldsymbol{\phi}}_{m}^{\alpha}\cdot{\boldsymbol{\phi}}_{n}^{\alpha})^{2}$
$+\frac{c^{\prime\prime\prime}}{4}\sum_{\alpha}\sum_{m<n}\sum_{m^{\prime}<n^{\prime}}({\boldsymbol{\phi}}^{\alpha}_{m}\cdot{\boldsymbol{\phi}}^{\alpha}_{n})({\boldsymbol{\phi}}^{\alpha}_{m^{\prime}}\cdot{\boldsymbol{\phi}}^{\alpha}_{n^{\prime}}).$ (6)
Note that $b$ terms are absent owing to the wavnumber conservation and $(m,n,m^{\prime},n^{\prime})=(1,2,3,4)$, $(1,3,2,4)$, and $(1,4,2,3)$ in the last term in the right-hand side. The coupling constants are denoted as
$a_{\rm L}$ $\equiv T-6J_{4},\quad J_{\rm L}\equiv-2J_{1}+6J_{3},$ (7)
$c^{\prime}=2c,\quad c^{\prime\prime}=4c,\quad c^{\prime\prime\prime}=8c.$ (8)
Note that ${\bm{k}}_{n}\equiv-{\bm{k}}_{n}$ and thus the ${\boldsymbol{\phi}}_{n}\cdot{\boldsymbol{\phi}}_{n}$ pair in the summation in $\sum_{{\bm{k}}}$ appears once. Since each of the fourth order terms are allowed by the cubic symmetry, we denote them as independent parameters.
For discussing a single-$\bm{q}$ order at the L points, $c^{\prime}$, $c^{\prime\prime}$, and $c^{\prime\prime\prime}$ terms are irrelevant and the fourth-order free energy is simply given by
$F_{L}^{1(4)}=\frac{c}{4}\sum_{\alpha}|{\boldsymbol{\phi}}^{\alpha}|^{4}.$ (9)
When multiple-$\bm{q}$ states are considered, the coupling constants $c^{\prime}$, $c^{\prime\prime}$, and $c^{\prime\prime\prime}$ matter as will be discussed in Sec. V.
II.4 $\Gamma$ point free energy
Under finite magnetic fields, the uniform quadrupole moment ${\boldsymbol{\phi}}_{0}^{\alpha}\equiv{\boldsymbol{\phi}}_{{\bm{k}}=\bf 0}^{\alpha}$ is induced via the effective field $\tilde{\bm{h}}$ in the second order perturbation of the magnetic field $\bm{H}$. The free energy up to the fourth-order terms for the $\Gamma$ point modes are
$F_{\Gamma}$ $=\sum_{\alpha}\Big{(}-\tilde{\bm{h}}\cdot{\boldsymbol{\phi}}_{0}^{\alpha}+\frac{a_{\Gamma 0}}{2}|{\boldsymbol{\phi}}_{0}^{\alpha}|^{2}\Big{)}+J_{\Gamma 0}{\boldsymbol{\phi}}_{0}^{\rm A}\cdot{\boldsymbol{\phi}}_{0}^{\rm B},$
$+\sum_{\alpha}\Big{\{}-bu_{0}^{\alpha}\Big{[}\big{(}u_{0}^{\alpha}\big{)}^{2}-3\big{(}v_{0}^{\alpha}\big{)}^{2}\Big{]}+\frac{c}{4}|{\boldsymbol{\phi}}_{0}^{\alpha}|^{4}\Big{\}},$ (10)
where
$a_{\Gamma 0}$ $=T+12J_{2}+6J_{4},\quad J_{\Gamma 0}=4J_{1}+12J_{3},$ (11)
and the effective field is given by
$\tilde{\bm{h}}=\begin{bmatrix}\tilde{h}_{u}\\ \tilde{h}_{v}\end{bmatrix}\equiv g\begin{bmatrix}2H_{z}^{2}-H_{x}^{2}-H_{y}^{2}\\ \sqrt{3}(H_{x}^{2}-H_{y}^{2})\end{bmatrix},$ (12)
and $g=(g_{J}\mu_{\rm B})^{2}[7/(3E_{4})-1/E_{5}]$ given by the excited $\Gamma_{4,5}$ CEF energy $E_{4,5}$.Hattori and Tsunetsugu (2014) $\mu_{\rm B}$ is the Bohr magneton and $g_{J}$ is the Lande’s $g$ factor for the multiplet with the angular momentum $J$; $g_{J}=4/5$ for the PrIr2Zn20 with $J=4$.
II.5 Mode coupling terms
We now discuss mode coupling terms consisting of the fields at the L points and those at $\Gamma$ points. The lowest order contributions come from the cubic terms as
$\delta F_{3}=$ $-3\lambda_{b}\sum_{\alpha,n}\Big{\{}u_{0}^{\alpha}\Big{[}\big{(}u_{n}^{\alpha}\big{)}^{2}-\big{(}v_{n}^{\alpha}\big{)}^{2}\Big{]}-v_{0}^{\alpha}(2u_{n}^{\alpha}v_{n}^{\alpha})\Big{\}}$
$=-3\lambda_{b}\sum_{\alpha,n}\phi_{0}^{\alpha}\big{(}\phi_{n}^{\alpha}\big{)}^{2}\cos(\theta^{\alpha}_{0}+2\theta^{\alpha}_{n}),$ (13)
Here we have parameterized
${\boldsymbol{\phi}}^{\alpha}_{\bf 0}$ $=\phi^{\alpha}_{0}\begin{bmatrix}\cos\theta^{\alpha}_{0}\\ \sin\theta^{\alpha}_{0}\end{bmatrix},\quad{\boldsymbol{\phi}}^{\alpha}_{n}=\phi^{\alpha}_{n}\begin{bmatrix}\cos\theta^{\alpha}_{n}\\ \sin\theta^{\alpha}_{n}\end{bmatrix},$ (14)
and the coupling constant $\lambda_{b}$ as an independent parameter. If $\lambda_{b}$ is estimated by Eq. (5), $\lambda_{b}=b$. The next leading order terms are the fourth order and given by
$\delta F_{4}=\sum_{\alpha,n}\left[\frac{\lambda^{(1)}_{c}}{2}|{\boldsymbol{\phi}}_{n}^{\alpha}|^{2}|{\boldsymbol{\phi}}_{0}^{\alpha}|^{2}+\lambda_{c}^{(2)}({\boldsymbol{\phi}}_{n}^{\alpha}\cdot{\boldsymbol{\phi}}_{0}^{\alpha})^{2}\right].$ (15)
Again, the coupling constant $\lambda_{c}^{(1,2)}$ has been introduced as an independent parameter. If $\lambda_{c}^{(1,2)}$ is estimated by Eq. (5), $\lambda^{(1)}_{c}=\lambda_{c}^{(2)}=c$.
In principle, ${\boldsymbol{\phi}}_{0}^{\alpha}$ can couple with ${\boldsymbol{\phi}}_{{\bm{k}}}^{\alpha}$ in the form of $\sim{\boldsymbol{\phi}}_{0}^{\alpha}{\boldsymbol{\phi}}_{\bm{k}}^{\alpha}{\boldsymbol{\phi}}_{-{\bm{k}}}^{\alpha}$, ${\boldsymbol{\phi}}_{0}^{\alpha}{\boldsymbol{\phi}}_{0}^{\alpha}{\boldsymbol{\phi}}_{\bm{k}}^{\alpha}{\boldsymbol{\phi}}_{-{\bm{k}}}^{\alpha}$. In this paper, we retain the order parameter ${\boldsymbol{\phi}}_{n}^{\alpha}$ and the induced moment ${\boldsymbol{\phi}}_{0}^{\alpha}$ under magnetic fields since the other fields are not important for the phase transition. Some renormalization effects arising from the field at the X points will be discussed in Sec. IV and Appendix B.
III Results: single-q orders
In this section, we carry out minimization of the free energy $F_{L}+F_{\Gamma}+\delta F_{3}+\delta F_{4}$ [Eqs. (6), (10), (13), and (13)] with the single-$\bm{q}$ assumptions. The purpose here is to understand how mode coupling effects affect the order parameter configuration. The results in this section will become bases for understanding the full phase diagram shown in Sec. IV where we sill take into account multiple-$\bm{q}$ configurations. Since we are interested in the phase transition owing to order parameter at the L point, we assume the second-order coefficients of the free energy at the other wavenumber are always positive. We first discuss the candidates of the stable configurations by analyzing the third and fourth order terms in $F_{\rm L}$. Then, we will show the numerical results of full temperature-field phase diagram.
III.1 Qualitative analysis
First, we notice that the inter-sublattice correlations are governed by $J_{\rm L}=-2J_{1}+6J_{3}$ for ${\boldsymbol{\phi}}_{n}^{\alpha}$ in Eq. (6). To reproduce the quadrupole configuration propposed by the neutron scattering experimentsIwasa et al. (2017), one needs $J_{\rm L}<0$. Thus, it is legitimate to set
${\boldsymbol{\phi}}_{\rm L}\equiv{\boldsymbol{\phi}}_{n}^{A}={\boldsymbol{\phi}}_{n}^{B}=\begin{bmatrix}u_{\rm L}\\ v_{\rm L}\end{bmatrix}=\phi_{\rm L}\begin{bmatrix}\cos\theta_{\rm L}\\ \sin\theta_{\rm L}\end{bmatrix}.$ (16)
Here, we have chosen one of the L points at ${\bm{k}}_{n}$ as a representative one since the four L points are equivalent even when $\bm{h}$ is finite. For the modes at the $\Gamma$ point, which are gapped and just induced by magnetic fields or by the square of the order parameter at the L point, we set
${\boldsymbol{\phi}}_{0}\equiv{\boldsymbol{\phi}}_{0}^{A}={\boldsymbol{\phi}}_{0}^{B}=\begin{bmatrix}u_{0}\\ v_{0}\end{bmatrix}=\phi_{0}\begin{bmatrix}\cos\theta_{0}\\ \sin\theta_{0}\end{bmatrix}.$ (17)
Second, we note that the single-$\bm{q}$ free energy $F_{L}$ for the L point order parameter is isotropic. This means there is no anisotropy at the transition temperature $T_{c}=6J_{4}+2|J_{1}-3J_{3}|$ at ${\bm{H}}={\bf 0}$. This isotropic nature, however, is destroyed by the mode coupling to the fields at the $\Gamma$ points through $\delta F_{3}$ and $\delta F_{4}$. When $T\lesssim T_{c}$, the induced ${\boldsymbol{\phi}}_{0}$ is sufficiently small, and thus the third and fourth order terms in $F_{\Gamma}$ can be irrelevant. In this situation the relative directions between ${\boldsymbol{\phi}}_{n}$ and ${\boldsymbol{\phi}}_{0}$ is determined by $\delta F_{3}$ and $\delta F_{4}$ as
$\cos(\theta_{0}+2\theta_{\rm L})=1\ {\rm and}\ \cos(\theta_{0}-\theta_{\rm L})=0.$ (18)
This leads to
$\theta_{0}=\frac{(2j-1)\pi}{3},\quad\theta_{\rm L}=\frac{(2j-1)\pi}{3}\pm\frac{\pi}{2},$ (19)
with $j=1,2,3$. These six domains are degenerate at ${\bm{H}}={\bf 0}\ (\tilde{\bm{h}}={\bf 0})$. See Fig. 2(a). We should comment that the three directions of ${\boldsymbol{\phi}}_{0}$ do not coincide with those favored by the $b(>0)$ term in $F_{\Gamma}:-b\phi_{0}^{3}\cos(3\theta_{0})$. Since ${\boldsymbol{\phi}}_{0}$ is not the primary order parameter, the third-order term is expected to play a minor role and the configurations (19) are robust at least $\tilde{\bm{h}}={\bf 0}$.
Figure 2: Schematic single-$\bm{q}$ order parameter configurations for (a) ${\bm{H}=0}$, (b) small ${\bm{H}}$ and low $T$, (c) small ${\bm{H}}$ and high $T$, and (d) large ${\bm{H}}$. (e) large ${\bm{H}}$. Blue (red) arrows indicate ${\boldsymbol{\phi}}_{0}$(${\boldsymbol{\phi}}_{\rm L}$). Filled red circles represent the directions $\theta_{L}$ given by Eq. (19) and open blue circles do $\theta_{\rm 0}$. A configuration for large ${\bm{H}}$ in the phase II in (c) is shown in (e) and that for ${\bm{H}}\parallel[110]$ in (f).
Now let us consider the cases for finite magnetic fields. For high symmetric directions $\bm{H}=\tfrac{H}{\sqrt{2}}(\sin\theta,\sin\theta,\sqrt{2}\cos\theta)$ with $20\leq\theta\leq\pi/$ 2, the effective field reads as $\tilde{h}_{u}=gH^{2}(3\cos^{2}\theta-1)$ and $\tilde{h}_{v}=0$. Note that $\tilde{\bm{h}}={\bf 0}$ at $\bm{H}\parallel[111]$ and $\tilde{h}_{u}>0$ for $\theta<\tan^{-1}\sqrt{2}\equiv\theta_{111}$, while $\tilde{h}_{u}<0$ for $\theta>\theta_{111}$. The effective field induces $u_{0}$, which leads to two “transition temperatures” $\tilde{T}_{c}^{u}$ for the $u_{\rm L}$ order and $\tilde{T}_{c}^{v}$ for $v_{\rm L}$. They are given by
$\tilde{T}_{c}^{u}$ $=T_{c}+6\lambda_{b}u_{0}-3\lambda_{c}u_{0}^{2},\ \tilde{T}_{c}^{v}=T_{c}-6\lambda_{b}u_{0}-\lambda_{c}u_{0}^{2},$ (20)
where we have set $\lambda_{c}^{(1)}=\lambda_{c}^{(2)}=\lambda_{c}$ for simplicity. Note that $\tilde{T}_{c}^{u,v}$ is a hypothetical transition temperature from the disordered state. The difference between the two is $\Delta\tilde{T}_{c}=\tilde{T}_{c}^{u}-\tilde{T}_{c}^{v}=2\lambda_{c}(\phi_{0}^{*}-u_{0})u_{0}$ with $\phi_{0}^{*}\equiv 6\lambda_{b}/\lambda_{c}$. For $u_{0}>0$, this means $\tilde{T}_{c}^{u}>\tilde{T}_{c}^{v}$ for $u_{0}<\phi_{0}^{*}$, i.e., for low fields, while $\tilde{T}_{c}^{v}>\tilde{T}_{c}^{u}$ for the higher field with $u_{0}>\phi_{0}^{*}$. In contrast, $\Delta\tilde{T}_{c}<0$ for $u_{0}<0$.
As is clear from the expression of $\Delta\tilde{T}_{c}$, it is expected that there is no additional phase transition within the single-$\bm{q}$ ${\boldsymbol{\phi}}_{n}$ sector for $u_{0}<0$, i.e., $\theta_{0}>\theta_{111}$. In the following, we will concentrate on the cases for $u_{0}>0$ ($\theta_{0}<\theta_{111}$). For $u_{0}>0$, the linear coupling between ${\boldsymbol{\phi}}_{0}$ and $\tilde{h}_{u}$ can alter the stable configuration from the above six domains (19) since their energy gain is given by the second order in ${\boldsymbol{\phi}}_{0}$. Since $\Delta\tilde{T}_{c}>0$, the phase transition from the disordered phase occurs in the $u_{\rm L}$ component: $\theta_{n}=0$ or $\pi$. Here, the two choices 00 and $\pi$ represent a simple translation and we will consider $\theta_{\rm L}=0$ below. For small $\tilde{h}_{u}$, $\delta F_{3}$ is more important than $\delta F_{4}$. This means $\theta_{0}=0$ satisfying $02\theta_{\rm L}+\theta_{0}=$ 0 [Fig. 2(b)]. Such a collinear state possesses the free energy $F_{\rm col}$ and the fluctuations around this is given as
$\delta F_{\rm col}$ $\simeq\lambda_{c}\phi_{0}\phi_{\rm L}^{2}\begin{bmatrix}\theta_{0}&\theta_{\rm L}\end{bmatrix}\hat{A}\begin{bmatrix}\theta_{0}\\ \theta_{\rm L}\end{bmatrix},$ (21)
$\hat{A}$ $=\begin{bmatrix}(\lambda_{c}\phi_{\rm L}^{2})^{-1}(\tilde{h}_{u}+9b\phi_{0}^{2})+\frac{1}{2}\phi_{0}^{*}-2\phi_{0}&\phi_{0}^{*}+2\phi_{0}\\ \phi_{0}^{*}+2\phi_{0}&2\phi_{0}^{*}-2\phi_{0}\end{bmatrix}.$ (22)
Indeed, the instability to finite $\theta_{{\rm L},0}$ occurs when one of the eigenvalue of the matrix $\hat{A}$ in Eq. (22) becomes zero. This takes place when
$\tilde{h}_{u}^{c1}=27\lambda_{b}\frac{\phi_{0}}{\phi_{0}^{*}-\phi_{0}}\phi_{\rm L}^{2}-9b\phi_{0}^{2}.$ (23)
Note that the factor $\phi_{0}^{*}-\phi_{0}$ of the denominator coincides with the factor in $\Delta T_{c}$. The configuration with $\theta_{0}\neq 0$ and $\theta_{\rm L}\neq 0$ gradually crossovers toward one for $\bm{H}=0$. See Fig. 2(c). When $\tilde{\bm{h}}$ is sufficiently large, $u_{0}$ grows and $\delta F_{4}$ plays an important role. Thus, ${\boldsymbol{\phi}}_{\rm L}$ is locked at $\theta_{\rm L}=\pm\pi/2$ as shown in Fig. 2(d). This is because $\delta F_{4}>\delta F_{3}$ for large $\tilde{h}_{u}$. By the similar analysis to Eqs. (21)-(23), the boundary determined by the field below which $|\theta_{\rm L}|<\pi/2$ [Fig. 2(e)] is given as
$\tilde{h}_{u}^{c2}=27\lambda_{b}\frac{\phi_{0}}{\phi_{0}-\phi_{0}^{*}}\phi_{\rm L}^{2}-9b\phi_{0}^{2}.$ (24)
For sufficiently large $\lambda_{b}$, the $\theta_{\rm L}=\pm\pi/2$ phase never appears because of $\tilde{T}_{c}^{v}<0$. Of course, for extremely large $\tilde{\bm{h}}$, the L point order vanishes trivially.
Finally, we briefly comment about the cases for $u_{0}<0$ ($\theta>\theta_{111}$) under finite magnetic fields. As we have noted, there is no phase transition once a quadrupole moment with $\theta_{\rm L}=\pm\pi/2$ is ordered as shown in Fig. 2(f). This is because $\theta_{\rm L}=\pm\pi/2$ and $\theta_{0}=\pi$ can minimize $\delta F_{3}$, $\delta F_{4}$, and $F_{\Gamma}$, simultaneously.
III.2 Phase diagram
Now we show the results of numerical minimization of the single-$\bm{q}$ free energy (6), (10), (13), and (15). We will show the result for ${\bm{H}}\parallel[001]$, i.e., $\tilde{h}_{u}>0$ and then for ${\bm{H}}\parallel[110]$, $\tilde{h}_{u}<0$. To understand the qualitative nature of the phase diagram, it is sufficient to discuss the cases for $J_{2}=J_{3}=0$.
Figures 3(a)–3(c) show the $T$-$H$ phase diagram for $\bm{H}\parallel[001]$, where the color represents the angle for the L point moment $\theta_{\rm L}$: (a) $b=0.1$, (b) $b=0.05$ , and (c) $b=0.01$ with $c=0.5$ fixed. As discussed in the previous section, the phase III is stabilized under high magnetic fields. This phase, however, is suppressed when $b$ becomes larger. We note that the collinear phase (phase I) appears for high $T$ and intermediate $H$. The reason why the phase I cannot appear at the lower $T$ is understood by seeing Eq. (23); there is no solution of the critical field $\tilde{h}_{u}^{c1}$ when $\phi_{\rm L}^{2}$, which is approximately $H$ independent, is large.
Figure 3: $T$$H$ phase diagram for $\bm{H}\parallel[001]$, $J_{1}=1.0$, $J_{2}=J_{3}=0.0$, $J_{4}=1.0$, and $c=0.5$. (a) $b=0.1$, (b) $b=0.05$, and (c) $b=0.01$. Color map represents the L point order parameter angle $\theta_{\rm L}$.
For $\bm{H}\parallel[110]$, the situation becomes quite simpler. Figure 4(a) shows the $T$-$H$ phase diagram for $\bm{H}\parallel[110]$. As discussed in the last part in the previous section, there is only one stable phase (phase IV), where $\theta_{\rm L}=\pm\pi/2$ and $\theta_{0}=\pi$.
Figures 4(b) and 4(c) show the real space quadrupole configurations with the ordering vector $\bm{k}_{1}$ for $\bm{H}\parallel[001]$ and $\bm{H}\parallel[110]$, respectively. For both cases, the quadrupole orders with $\bm{\phi}_{A}=\bm{\phi}_{B}$ are ferroic on a (111) plane, while antiferroic between the neighboring (111) plane as shown in different colors in Fig. 4. When the ordering vector is $\bm{k}_{n}(n=2,3,4)$, the quadrupole configurations are ferroic on the plane perpendicular to ${\bm{k}}_{n}$ with $\bm{\phi}_{A}=-\bm{\phi}_{B}$.
Figure 4: (a) $T$$H$ phase diagram for $\bm{H}\parallel[110]$, $J_{1}=-1$, $J_{2}=J_{3}=0$, $J_{4}=1$, $c=0.5$, and $b=0.05$. Color map represents the L point order parameter angle $\theta_{\rm L}$. (b) $\theta$$H$ phase diagram from [001]($\theta=0$) to [110] ($\theta=\pi/2$).
IV Results: multiple-q orders
In this section, we consider multiple-$\bm{q}$ orders for the quadrupole moments at the L points. We will first introduce a simple symmetric multiple-$\bm{q}$ states with the same amplitudes for all the $\bm{q}_{n}$ in Sec. IV.1. Then, we will show numerical results, where we relax the constraint about the amplitudes of the order parameters.
IV.1 Free energy for symmetric states
In this section, we consider double-$\bm{q}\ (\mathcal{N}=2)$ and quadruple-$\bm{q}\ (\mathcal{N}=4)$ configurations, whose real space moments are given by
${\boldsymbol{\phi}}^{\alpha}(\bm{r})=\begin{bmatrix}u^{\alpha}_{\bm{r}}\\ v^{\alpha}_{\bm{r}}\end{bmatrix}=\frac{\Phi_{\alpha}}{\sqrt{\mathcal{N}}}\sum_{n=1}^{\mathcal{N}}e^{i{\bm{k}}_{n}\cdot\bm{r}}\begin{bmatrix}\cos\theta_{n}^{\alpha}\\ \sin\theta_{n}^{\alpha}\\ \end{bmatrix},$ (25)
where $\bm{r}$ is labeled by the A sublattice position as before and we have assumed that the magnitude for each ${\bm{k}}_{n}$ is the same for simplicity (symmetric states). In this normalization $(\mathcal{N})$, the quadratic part is unchanged from the single-$\bm{q}$ expression. Thus, types of order realized below the second-order transition at $T=T_{c}$ are determined by the fourth-order terms.
For double-$\bm{q}$ states, the fourth-order terms are
$F_{\rm L}^{2(4)}=\sum_{\alpha}\Phi_{\alpha}^{4}\left[\frac{c}{8}+\frac{c^{\prime}}{16}+\frac{c^{\prime\prime}}{16}(c^{\alpha}_{12})^{2}\right].$ (26)
Here, $c^{\alpha}_{mn}\equiv\cos(\theta^{\alpha}_{m}-\theta^{\alpha}_{n})$. This clearly indicates that $\theta_{1}^{\alpha}-\theta_{2}^{\alpha}=\pm\pi/2$ for $c^{\prime\prime}>0$ (orthogonal double-$\bm{q}$), while $\theta_{1}^{\alpha}-\theta_{2}^{\alpha}=0$ or $\pi$ for $c^{\prime\prime}<0$ (parallel double-$\bm{q}$). As explained in Appendix A, the double-$\bm{q}$ configurations consist of four ferroic chains running along the direction of ${\bm{k}}_{3}+{\bm{k}}_{4}$ when only the A-sublattice sites are concerned.
For quadruple-$\bm{q}$ states,
$F_{\rm L}^{4(4)}=$ $\sum_{\alpha}\Phi_{\alpha}^{4}\Bigg{[}\frac{c}{16}+\frac{3c^{\prime}}{32}+\frac{c^{\prime\prime}}{64}\sum_{m<n}(c^{\alpha}_{mn})^{2}$
$+$ $\frac{c^{\prime\prime\prime}}{64}(c_{12}^{\alpha}c_{34}^{\alpha}+c_{13}^{\alpha}c_{24}^{\alpha}+c_{14}^{\alpha}c_{23}^{\alpha})\Bigg{]}.$ (27)
Note that the $c^{\prime\prime\prime}$ term can be negative, e.g., when $\theta_{1}=\theta_{2}=\theta_{3}=\theta_{4}+\pi$. This means that when $c^{\prime\prime\prime}$ is large, such quadruple-$\bm{q}$ orders are favored. The lowest free energy under the constraint (8) is $F_{\rm L}^{4(4)}=\tfrac{c}{4}\sum_{\alpha}\Phi_{\alpha}^{4}$ with $(\theta_{1},\theta_{2},\theta_{3},\theta_{4})=(\theta,\theta,\theta^{\prime},\theta^{\prime}+\pi)$, where $\theta$ and $\theta^{\prime}$ are arbitrary and those with any two $\theta_{n}$’s permutations are equivalent ones. The unit cell for the quadrupole-$\bm{q}$ configurations contains eight A-sublattice sites and thus sixteen sites in total. See also Fig. 9 in Appendix A.
In Eqs. (26) and (27), $\theta_{n}$ dependence appears only in $c^{\prime\prime}$ and $c^{\prime\prime\prime}$ terms. Figure 6 shows the phase diagram at $T=T_{c}$ on the $c^{\prime\prime\prime}$$c^{\prime\prime}$ plane for (a) $c^{\prime}>2c$ and (b) $c^{\prime}<2c$ with $c>0$. For large $c^{\prime}$, multiple-$\bm{q}$ states are not realized as is evident from the expressions (26) and (27). See Fig. 6(a). The conditions given by Eq. (8) just correspond to the phase boundary for $c^{\prime\prime},c^{\prime\prime\prime}>0$ and $c^{\prime}=2c$, where all the single-$\bm{q}$, double-$\bm{q}$, and quadruple-$\bm{q}$ are degenerate. Note that away from the phase boundaries between quadruple-$\bm{q}$ and double- or single-$\bm{q}$ phases, the direction of the $\bm{\phi}_{n}$ in the quadruple-$\bm{q}$ states are pinned to $(\theta_{1}^{\alpha},\theta_{2}^{\alpha},\theta_{3}^{\alpha},\theta_{4}^{\alpha})$$=$(0,0,0,0)$0$(0,0,0,0)( 0 , 0 , 0 , 0 ), $(0,0,\pi,\pi)$, $(0,\pi,\pi,0)$, $(0,\pi,0,\pi)$ for $c^{\prime\prime\prime}<0$, while $(\theta_{1}^{\alpha},\theta_{2}^{\alpha},\theta_{3}^{\alpha},\theta_{4}^{\alpha})$$=$$(0,0,0,\pi)$, $(0,0,\pi,0)$, $(0,\pi,0,0)$, $(\pi,0,0,0)$ for $c^{\prime\prime\prime}>0$. Here, global sign changes $\bm{\phi}^{\alpha}\to-\bm{\phi}^{\alpha}$: $\theta_{n}^{\alpha}\to\theta_{n}^{\alpha}+\pi$, have not been included since they are also realized by trivial translations.
Figure 5: Phase diagram at $T=T_{c}$ in the $c^{\prime\prime\prime}$$c^{\prime\prime}$ plane. (a) $c^{\prime}>2c$, where the phase boundary between the single-$\bm{q}$(1q) and quadruple-$\bm{q}$(4q) is given by $c^{\prime\prime}=|c^{\prime\prime\prime}|/2-(c^{\prime}-2c)$. (b) $c^{\prime}<2c$, where the phase boundary between the double-$\bm{q}$(2q) and quadruple-$\bm{q}$(4q) is given by $c^{\prime\prime}=|c^{\prime\prime\prime}|/2+(2c-c^{\prime})/3$.
So far, we have treated the fourth-order coefficients $c$’s as a phenomenological parameters. Here, we discuss the renormalization owing to modes other than $\bm{\phi}_{0}$ and $\bm{\phi}_{n}$. In the lowest orders, the cubic terms in Eq. (5) are the most important. Among them, there are special terms which connect two L-point modes at $\bm{q}_{m}$ and $\bm{q}_{n}$ ($m\neq n$) via $\sim\phi_{\bm{k}}\phi_{m}\phi_{n}$, where ${\bm{k}}$ is one of the wavenumbers at the X points $\bm{q}_{\rm X}=(2\pi,0,0),\bm{q}_{\rm Y}=(0,2\pi,0)$, or $\bm{q}_{\rm Z}=(0,0,2\pi)$. The combination of $({\bm{k}},\bm{q}_{m},\bm{q}_{n})$ is restricted by the Umklapp condition ${\bm{k}}+\bm{q}_{m}+\bm{q}_{n}$ and there are two sets of $(m,n)$ for a given ${\bm{k}}$. As shown in Appendix B, these couplings lead to suppression of the coefficients $c^{\prime}$ and $c^{\prime\prime\prime}$. See Eq. (38). Furthermore, nontrivial couplings between the $\Gamma$ and the L points fields appear after integrating out the X point fields as $\sim\phi_{0}\phi_{m}^{2}\phi_{n}^{2}$ and $\sim\phi_{0}\phi_{1}\phi_{2}\phi_{3}\phi_{4}$. They play an important role under magnetic fields $\bm{H}$ since $\phi_{0}$ is induced by $\bm{H}$, and thus, these terms renormalize $c$’s with varying their magnitude depending on $\bm{H}$.
The suppression of $c^{\prime}$ and $c^{\prime\prime\prime}$ due to the renormalization with retaining $c^{\prime\prime}>0$ suggests that the orthogonal double-$\bm{q}$ states are more stable. However, the coupling between the fields at the $\Gamma$ and the L points possesses an energy gain for the single-$\bm{q}$ orders and double-$\bm{q}$ ones with $\phi_{m}\neq\phi_{n}$. This is because there is a finite $\phi_{0}$ induced in the single-$\bm{q}$ states, while there is no gain for the double-$\bm{q}$ states with $\phi_{n}=\phi_{m}$ with $\theta_{n}-\theta_{m}=\pm\pi/2$. See the coupling in Eq. (13). As we will see in Sec. IV.2, the nontrivial terms generated by the integration of the X points fields choose some particular $\theta$ among degenerate orthogonal double-$\bm{q}$ configurations.
IV.2 Phase diagram
Now, we discuss $T$$H$ phase diagram ($H\equiv|\bm{H}|$) taking into account multiple-$\bm{q}$ orders in addition to the single-$\bm{q}$ ones. We numerically minimize the free energy $F_{\rm L}+F_{\Gamma}+\delta F_{3}+\delta F_{4}+\delta F$, where $\delta F$ is due to the integration of the X point fields as discussed in Appendix B and the others are given in Eqs. (6), (10), (13), and (15). Figures 6 and 8 show phase diagrams for $\bm{H}\parallel[001]$, i.e., $\tilde{\bm{h}}=\tilde{h}_{u}(1,0)$ $(\tilde{h}_{u}\geq 0)$, and double-$\bm{q}$ configurations in the quadrupole space and real space. Figure 6(a) shows the phase diagram for $J_{1}=1$, $c=0.5$, and $b=0.1$. The exchange coupling constants are the same as used in Fig. 3: $(J_{2},J_{3},J_{4})=(0.0,0.0,1.0)$. Except for the low-field region, the phase diagram is similar to those shown in Fig. 3, showing the single-$\bm{q}$ orders I, II, and III. An orthogonal double-$\bm{q}$ configuration I2q with $\bm{\phi}_{{\bm{k}}_{1}}=(\phi_{1},0)$ and $\bm{\phi}_{{\bm{k}}_{2}}=(0,\phi_{2})$ is stabilized for finite magnetic fields at high $T$ region below the phase I. Here, we take a domain with the ordering vectors ${\bm{k}}_{1}$ and ${\bm{k}}_{2}$. The configuration of I2q state in the quadrupole space is shown in Fig. 8(a) and its real-space one is illustrated in Fig. 8(b). At ${\bm{H}}=0$, the orthogonal double-$\bm{q}$ states with $\phi_{1}=\phi_{2}$ discussed in Sec. IV.1 are realized and degenerate in the angle $\theta=0$. As ${\bm{H}}$ increases, $\phi_{1}$ becomes larger than $\phi_{2}$. Since the phase I is the single-$\bm{q}$ order with $v_{\rm L}=0$, the symmetry breaking from the phase I to the phase I2q is caused by the emergence of $\bm{\phi}_{2}=(0,\phi_{2})$ through the second-order transition. In this sense, $\phi_{1}>\phi_{2}$ is quite natural. In the free energy, this is due to the $\delta F_{3}$, which is minimized when the L point field is parallel to $\bm{\phi}_{0}=(u_{0},0)$, while maximized when it is perpendicular. As for the phase boundary between the phase I2q and the phase II, it is of first-order, while the other phase boundaries are second-order.
At lower $T$, the induced $\Gamma$ point moment $\bm{\phi}_{0}$ favors the configuration III2q shown in Figs. 8(c) and 8(f), where $\bm{\phi}_{1}=\tfrac{1}{\sqrt{2}}\phi_{1}(1,1)$ and $\bm{\phi}_{2}=\tfrac{1}{\sqrt{2}}\phi_{2}(-1,1)$ with $\phi_{1}=\phi_{2}$. This configuration is stabilized by the nontrivial coupling discussed in the last part of Sec. IV.1. See also the last term in Eq. (39). In Fig. 6(a), an intermediate phase II2q toward the phase III2q appears at low temperatures. The quadrupole configurations in the phase II2q are “rotated” from those in the phase I2q with keeping $\bm{\phi}_{1}\perp\bm{\phi}_{2}$ as shown in Fig. 8(b). For smaller $b=0.05$ [Fig. 6(b)], the phase I, II, and the double-$\bm{q}$ orders shrink, while the phase III expands. The overall phase diagram, however, is unchanged.
As an example for more general cases, we show the phase diagram for $(b,J_{2},J_{3},J_{4})=(0.03,0.5,0.3,0.4)$ in Fig. 6(c). For this parameter set, the second-order coefficient $a_{\rm X}$ of the X point fields almost vanish at $T=0$ as shown in Fig. 1(c), the renormalization effects due to the X point fields grows rapidly at low temperatures. This causes the strong stabilization of the phase III2q in Fig. 6(c), showing reentrant behavior as a function of $H$. Interestingly, there is a thin region of the phase II2q. The phase boundary between the phase II2q and II is of the second order. Similarly, that between the small region of phase I2q and III is also second order. We note that the boundaries indicated by the dashed lines are not symmetry breaking phase boundary since the symmetry of the three double-$\bm{q}$ phases are the same as is evident from the real space configurations in Figs. 8(d)–8(f). See also the discussion in Appendix C.
For $\bm{H}\parallel[110]$, a phase I2q appears at low field region as shown in Fig. 7. In contrast to the phase I2q for $\bm{H}\parallel[001]$, the quadrupole configuration is $\bm{\phi}_{1}=(0,\phi_{1})$ and $\bm{\phi}_{2}=(\phi_{2},0)$ with $\phi_{1}>\phi_{2}$. This is natural since the neighboring single-$\bm{q}$ phase is the phase IV. The sharp increase of the critical field between I2q and IV phases are due to the fact that $a_{\rm X}$ is close to 0 in (c).
Figure 6: $T$$H$ phase diagram for $\bm{H}\parallel[001]$, $J_{1}=1.0$, and $c=0.5$. (a) $(b,J_{2},J_{3},J_{4})=(0.1,0.0,0.0,1.0)$, (b) $(0.05,0.0,0.0,1.0)$, and (c) $(0.03,0.5,0.3,0.4)$. Color represents $\Delta\phi\equiv|\phi_{1}-\phi_{2}|$, where double-$\bm{q}$ orders with the ordering vectors ${\bm{k}}_{1}$ and ${\bm{k}}_{2}$ are assumed. For single-$\bm{q}$ orders, we use the convention of $\phi_{2}=0$. Thick lines represent phase boundaries, where “1st” represents a line of first-order phase transition. Dashed lines are boundaries among double-$\bm{q}$ configurations. Low $T$ regime is not shown since the Landau expansion is not valid in general.
Figure 7: $T$$H$ phase diagram for $\bm{H}\parallel[001]$, $J_{1}=1.0$, and $c=0.5$. (a) $(b,J_{2},J_{3},J_{4})=(0.1,0.0,0.0,1.0)$, (b) $(0.05,0.0,0.0,1.0)$, and (c) $(0.03,0.5,0.3,0.4)$, which correspond to those in Fig. 7(a), 7(b), and 7(c), respectively. Color represents $\Delta\phi\equiv|\phi_{1}-\phi_{2}|$, similarly to Fig. 6
Figure 8: Schematic orthogonal double-$\bm{q}$ order parameter configurations for (a) I2q: $(\theta_{1},\theta_{2})=(0,\pi/2)$ with $\phi_{1}>\phi_{2}$, (b) II2q: $(\theta_{1},\theta_{2})=(\theta,\theta+\pi/2)$ with $\phi_{1}>\phi_{2}$, and (c) III2q: $(\theta_{1},\theta_{2})=(\pi/4,3\pi/4)$ and $\phi_{1}=\phi_{2}$. Arrows indicate the order parameters $\bm{\phi}_{1}$ (L1), $\bm{\phi}_{2}$ (L2), and $\bm{\phi}_{0}$ $(\Gamma)$. Four small circles represent quadrupole moments at the A sublattice real-space positions $\bm{r}=(0,0,0),(1,0,0),(\tfrac{1}{2},\frac{1}{2},0)$, and $(\tfrac{1}{2},\frac{1}{2},1)$. Real space L point quadrupole configurations for (d) $(\theta_{1},\theta_{2})=(0,\pi/2)$ and $\phi_{1}=\sqrt{3}\phi_{2}$ in the phase I2q, (e) $(\theta_{1},\theta_{2})=(\pi/24,13\pi/24)$ and $\phi_{1}=1.2\phi_{2}$ in the phase II2q, and (f) $(\theta_{1},\theta_{2})=(\pi/4,3\pi/4)$, and $\phi_{1}=\phi_{2}$ in the phase III2q. Note that only L point quadrupole components without those at the $\Gamma$ point are shown. Skyblue and gray arrows represent the directions of ${\bm{k}}_{1}$ and $-{\bm{k}}_{2}$. For guide to eyes, some of the nearest-neighbor bonds are connected by white bonds. Circles are the A sublattice atoms in the diamond structure and their colors represent the sublattice (red, green, blue, white). For clarity, the B sublattice atoms are not drawn. Colored bonds running [01¯$1]$ 1 ] directions connects the A sublattice sites with identical quadrupole moments.
V Discussions
V.1 Comparison with the experimental data of PrIr2Zn20
PrIr2Zn20 shows several phase transitions under magnetic field at low temperatures as reported in the early study.Onimaru et al. (2011) However, the order parameters have not been identified except for the $x^{2}-y^{2}$ type quadrupole order with the ordering vector at the L point for $\bm{H}\parallel[110]$ by the neutron scattering experimentIwasa et al. (2017). The analysis of this study is fully based on this fact. Namely, we have assumed that the leading order parameter is the quadrupole moment at the L points. Constructing the symmetry-allowed free energy consisting of them and the related degrees of freedom, we have analyzed the phase diagram for $\bm{H}\parallel[001]$ in addition to [110]. Our analysis demonstrates that a $2z^{2}-x^{2}-y^{2}$ configuration with the ordering vector at the L point is stable at high-$T$ and up to intermediate magnetic field strength for $\bm{H}\parallel[001]$. This suggests that the $2z^{2}-x^{2}-y^{2}$ order can be a promising candidate for the high-$T$ phase found for $\bm{H}\parallel[001]$ in the experimental worksOnimaru et al. (2011); Kittaka et al. (2024). As for the order parameter at zero field, the single-$\bm{q}$ configuration is not stable while the double-$\bm{q}$ I2q is realized with $\phi_{1}=\phi_{2}$. For finite ${\bm{H}}\parallel[001]$, this double-$\bm{q}$ configuration is modified and $\bm{\phi}_{1}=(\phi_{1},0)$ and $\bm{\phi}_{2}=(0,\phi_{2})$ with $\phi_{1}>\phi_{2}$ is favored. $\phi_{2}$ increases as lowering $T$ and there appear phase II2q and III2q by simultaneously rotating $\bm{\phi}_{1}$ and $\bm{\phi}_{2}$. As the cubic anisotropy $b$ increases, the influence of the renormalized fourth-order coefficients becomes larger and favors the phase II2q and III2q.
As magnetic field increases, the double-$\bm{q}$ order is destabilized and single-$\bm{q}$ orders take place as discussed in Sec. III. Thus, as increasing $\bm{H}\parallel[001]$, the double-$\bm{q}$, the single-$\bm{q}$ phase II with both $u_{\rm L}$ and $v_{\rm L}$ finite, and then the phase III with only $v_{\rm L}$ are realized. The trace of the phase boundaries between them have not been clearly observed in the experiments. However, there are several anomalies in the ultrasonic experimentsIshii et al. (2011). We hope the relation between our phase diagram and these anomalies can be clarified in future experimental studies. In addition to this, there is ambiguity about the stability of the phase III, whether the phase III extends toward the higher field as shown in Fig. 3(c) or occupies only small area [Fig. 3(a)]. This can be controlled by the cubic coupling $b$. For $\bm{H}\parallel[110]$, there is no high-T phase in contrast to the case for $\bm{H}\parallel[001]$ in the experiments.Onimaru et al. (2011); Kittaka et al. (2024) This point is consistent with our results. Other anomalies observed in the ultrasound experimentsIshii et al. (2011) similar to those for $\bm{H}\parallel[001]$ are also interesting to examine in more detail in the future works. Since the anomalies are tiny and thus the comparison between the experiments and the present theory is not straightforward, while there is a phase boundary between the phase I2q and IV in Fig. 7.
We also note that the phase diagram as a function of magnetic field from $\bm{H}\parallel[001]$ to $\bm{H}\parallel[110]$ is described in a function of angle $\Theta$, where $\bm{H}=\tfrac{1}{2}H(\sin\Theta,\sin\Theta,2\cos\Theta)$. In Eq. (12), the effective field to the quadrupole moments is written as $\tilde{\bm{h}}=gH^{2}(3\cos^{2}\Theta-1,0)$. Thus, the critical field of quadrupole orders for the angle $\Theta$, $H_{c}(\Theta)$ is expected to be
$H_{c}(\Theta)=\frac{\sqrt{2}H_{c}(0)}{\sqrt{3\cos^{2}\Theta-1}}.$ (28)
Such behavior can be one of strong evidences for quadrupole orders. We hope the future experimental studies on the phase diagram at finite fields clarify this point.
V.2 Future perspectives
We have shown that single- and double-$\bm{q}$ quadrupole orders are realized in our study as a phenomenological model for PrIr2Zn20. The orthogonal double-$\bm{q}$ orders are favored by the effective fourth-order terms due to the X point fields, while the quadruple ones are not. Although we are interested in the orbital orders at the L points in this paper, quadruple-$\bm{q}$ orders at incommensurate ordering vectors along [111] directions has been studied in cubic magnets, showing hedgehog type topological magnetic structure.Okumura et al. (2020, 2022) In principle, such exotic phases are also possible for the orbital model discussed in this paper when the exchange coupling favors incommensurate ordering vectors by introducing interaction frustration or anisotropic interaction ignored in this studyHattori et al. (2024).
For quantitative comparison between the theoretical and experimental results, it is also needed to incorporate the actual crystal electric field states in the real materials. They vary under magnetic fields and thus the parameters in the free energy can change. For theoretical or computational aspects, it is also interesting to take into account the fluctuation effects by using e.g., classical Monte Carlo simulation employed in the previous studies on Pr-based materials.Hattori and Tsunetsugu (2016); Freyer et al. (2018, 2020) The thermal fluctuation may alter the mean-field phase diagram and to check their stability is important issue as one of our future problems.
VI Summary
We have analyzed quadrupole orders with its modulation characterized by the wavenumber at the L points. Their Landau free energy is constructed under magnetic fields $\bm{H}$. We find that the phase diagrams show strong anisotropy. We have clarified that $x^{2}-y^{2}$ type quadrupole order is stabilized along $\bm{H}\parallel[110]$, which is consistent with that proposed in the neutron scattering experiment. We also propose that the unidentified high-temperature phase for $\bm{H}\parallel[001]$ in the experiments is a collinear $2z^{2}-x^{2}-y^{2}$ order and the orthogonal double-$\bm{q}$ configurations are stable at low fields. These results offers further experimental examination of the order parameters in PrIr2Zn20 and related compounds.
Acknowledgment
The authors thank T. Onimaru T. Ishitobi, and S. Kittaka for fruitful discussion. This work was supported by JSPS KAKENHI (Grant Nos. JP23K20824 and JP23H04869).
Appendix A Multiple-q configuration
In this Appendix, we show the real-space modulations for the double-$\bm{q}$ and quadruple-$\bm{q}$ orders since they are rather complicated. The real-space configurations for the double- and quadrupole-$\bm{q}$ orders are given by Eq. (25). Here, we concentrate on the sublattice independent orders $\bm{\phi}(\bm{r})$ with $\Phi^{\alpha}=\Phi$ and $\theta^{\alpha}_{n}=\theta_{n}$ since only such orders appear in this paper. Thus, one rewrites Eq. (25) as
$\bm{\phi}(\bm{r})=\mathbb{E}_{\mathcal{N}}(\bm{r})\cdot\begin{pmatrix}\bm{\phi}_{1}\\ \bm{\phi}_{2}\\ \cdots\end{pmatrix},\ \bm{\phi}_{n}=\frac{\Phi_{n}}{\sqrt{\mathcal{N}}}\begin{pmatrix}\cos\theta_{n}\\ \sin\theta_{n}\end{pmatrix},$ (29)
where $n=1,2,\cdots,\mathcal{N}$ and $\mathbb{E}_{\mathcal{N}}$ is an ${\mathcal{N}}$ dimensional vector
$\mathbb{E}_{2}(\bm{r})$ $=(e^{i{\bm{k}}_{1}\cdot\bm{r}},e^{i{\bm{k}}_{2}\cdot\bm{r}}),$ (30)
$\mathbb{E}_{4}(\bm{r})$ $=(e^{i{\bm{k}}_{1}\cdot\bm{r}},e^{i{\bm{k}}_{2}\cdot\bm{r}},e^{i{\bm{k}}_{3}\cdot\bm{r}},e^{i{\bm{k}}_{4}\cdot\bm{r}}).$ (31)
Here, we have assumed, without loss of generality, double-$\bm{q}$ orders with ${\bm{k}}_{1}$ and ${\bm{k}}_{2}$. Since $\bm{r}$ is the position in the fcc lattice, there are four choices for $\mathbb{E}_{2}(\bm{r}):(\pm 1,\pm 1)$, and eight choices for $\mathbb{E}_{4}(\bm{r}):\pm(1,1,1,1),\ \pm(1,1,-1,-1),\ \pm(1,-1,1,-1)$, and ±$\pm(-1,1,1,-1)$. In Fig. 9, $\mathbb{E}_{\mathcal{N}}(\bm{r})$’s are illustrated and the colors distinguish the vectors $\mathbb{E}_{2,4}$. Note that only the A-sublattice positions in the diamond structure are shown. The unit cell for the double-$\bm{q}$ orders contains four A sublattice sites, each of which forms ferroic chain along the direction parallel to ${\bm{k}}_{3}+{\bm{k}}_{4}\parallel(0,-1,1)$ as shown in Fig. 9(a). The primitive translation vector along this direction is unchanged, while those perpendicular to it are modified to $\frac{1}{2}(2,1,1)$ and $\frac{1}{2}(2,-1,-1)$. For quadruple-$\bm{q}$ orders, the primitive translational vectors are as twice large as those for the original fcc lattice: (1,1,0), (1,0,1), and (0,1,1). The unit cell contains eight sites for the quadruple-$\bm{q}$ orders and sixteen when the B-sublattice is included.
Figure 9: (a) Double-$\bm{q}$ configuration $\mathbb{E}_{2}$ at the A sublattice. Tubes connect the same $\mathbb{E}_{2}$’s. Thin lines indicates the NN bonds of diamond structure. Arrows indicates the direction of ${\bm{k}}_{1}$ and $-{\bm{k}}_{2}$. (b) Quadruple-$\bm{q}$ configuration $\mathbb{E}_{4}$. (c) Symmetric quadruple-$\bm{q}$ configuration for $(\theta_{1},\theta_{2},\theta_{3},\theta_{4})=(0,0,0,0)$ and (d) for $(\theta_{1},\theta_{2},\theta_{3},\theta_{4})=(0,0,\pi/2,3\pi/2)$. For (d), there are four kinds of ferroic chains. Two of them are similar to the double-$\bm{q}$ configuration for $x=2n$($n:$ integers) planes with the $2z^{2}-x^{2}-y^{2}$ orbital, while the other two chains for $x=2n+1$ plane with the $x^{2}-y^{2}$ orbital run along ${\bm{k}}_{1}+{\bm{k}}_{2}$ direction.
Since several double-$\bm{q}$ configurations are shown in Fig. 8, we here examine the real space configuration of the symmetric quadruple-$\bm{q}$ orders discussed in Sec. IV.1, although they are not stabilized in the $T$$H$ phase diagram in Fig. 6. It is evident that the configurations with “collinear” types such as $(\theta_{1},\theta_{2},\theta_{3},\theta_{4})=(0,0,0,0)$ and $(\theta_{1},\theta_{2},\theta_{3},\theta_{4})=(0,0,0,\pi)$ are partially ordered states. For example, the quadrupole moments vanish at the sites where $\sum_{n}e^{i{\bm{k}}_{n}\cdot\bm{r}}=0$ for the former as shown in Fig. 9(c). As an example of “non-collinear” case, Fig. 9(d) illustrates the configuration for $(\theta_{1},\theta_{2},\theta_{3},\theta_{4})=(0,0,\pi/2,3\pi/2)$.
Appendix B Renormalization effects due to the fields at the X points
In this Appendix, we consider the influence of the modes $\bm{Q}_{\ell}^{\alpha}=(Q_{\ell,z}^{\alpha},Q_{\ell,x}^{\alpha})$ ($\ell=$ X, Y, Z) at the X points: $\bm{q}_{\rm X}=(2\pi,0,0),\bm{q}_{\rm Y}=(0,2\pi,0)$, and $\bm{q}_{\rm Z}=(0,0,2\pi)$. They are more important than other modes at general wavenumbers since, for example, $\bm{q}_{1}+\bm{q}_{2}+\bm{q}_{\rm X}\equiv\bf{0}$ leads to cubic mode coupling with the two different fields at the L points.
The free energy for the X points fields are sublattice diagonal and $F_{\rm X}=\sum_{\alpha={\rm A,B}}F_{\rm X}^{\alpha}$, where
$F^{\alpha}_{\rm X}=$ $\frac{1}{2}\sum_{\ell={\rm X,Y,Z}}\bm{Q}^{\alpha}_{\ell}\begin{bmatrix}a_{\rm X}-6bu_{0}^{\alpha}&6bv_{0}^{\alpha}\\ 6bv_{0}^{\alpha}&a_{\rm X}+6bu_{0}^{\alpha}\end{bmatrix}\bm{Q}^{\alpha}_{\ell}$
$-6b\sum_{\ell={\rm X,Y,Z}}\sum_{i=1}^{2}\Big{[}Q_{\ell,z}^{\alpha}C_{\ell i}^{\alpha}-Q_{\ell,x}^{\alpha}S_{\ell i}^{\alpha}\Big{]}+\cdots.$ (32)
Here, the fourth-order terms and the cubic term for three X points fields are neglected since the modes at the X points are not critical. $a_{\rm X}$ and $C$’s and $S$’s consisting of two L points variables are given as
$a_{\rm X}\equiv T-4J_{2}+6J_{4},$ (33)
$\begin{split}&C_{{\rm X}1}^{\alpha}=u_{1}^{\alpha}u_{2}^{\alpha}-v_{1}^{\alpha}v_{2}^{\alpha},\quad C_{{\rm X}2}^{\alpha}=u_{3}^{\alpha}u_{4}^{\alpha}-v_{3}^{\alpha}v_{4}^{\alpha},\\ &S_{{\rm X}1}^{\alpha}=u_{1}^{\alpha}v_{2}^{\alpha}+v_{2}^{\alpha}u_{1}^{\alpha},\quad S_{{\rm X}2}^{\alpha}=u_{3}^{\alpha}v_{4}^{\alpha}+v_{3}^{\alpha}u_{4}^{\alpha},\\ &C_{{\rm Y}1}^{\alpha}=u_{1}^{\alpha}u_{4}^{\alpha}-v_{1}^{\alpha}v_{4}^{\alpha},\quad C_{{\rm Y}2}^{\alpha}=u_{2}^{\alpha}u_{3}^{\alpha}-v_{2}^{\alpha}v_{3}^{\alpha},\\ &S_{{\rm Y}1}^{\alpha}=u_{1}^{\alpha}v_{4}^{\alpha}+v_{1}^{\alpha}u_{4}^{\alpha},\quad S_{{\rm Y}2}^{\alpha}=u_{2}^{\alpha}v_{3}^{\alpha}+v_{2}^{\alpha}u_{3}^{\alpha},\\ &C_{{\rm Z}1}^{\alpha}=u_{1}^{\alpha}u_{3}^{\alpha}-v_{1}^{\alpha}v_{3}^{\alpha},\quad C_{{\rm Z}2}^{\alpha}=u_{2}^{\alpha}u_{4}^{\alpha}-v_{2}^{\alpha}v_{4}^{\alpha},\\ &S_{{\rm Z}1}^{\alpha}=u_{1}^{\alpha}v_{3}^{\alpha}+v_{1}^{\alpha}u_{3}^{\alpha},\quad S_{{\rm Z}2}^{\alpha}=u_{2}^{\alpha}v_{4}^{\alpha}+v_{2}^{\alpha}u_{4}^{\alpha}.\end{split}$ (34)
By minimizing $F_{\rm X}$ with respect to the X points fields, we obtain their stationary values as
$Q_{\ell,z}^{\alpha}=\frac{6b}{a_{\rm X}}\sum_{i=1}^{2}C^{\alpha}_{{\ell}i},\quad Q^{\alpha}_{\ell,x}=-\frac{6b}{a_{\rm X}}\sum_{i=1}^{2}S^{\alpha}_{{\ell}i}.$ (35)
Substituting them into Eq. (32), we obtain up to the linear order in $\phi_{0}$: $\delta F=\sum_{\alpha={\rm A,B}}\delta F^{\alpha}_{0}+\delta F^{\alpha}_{1}+\cdots$ as
$\delta F^{\alpha}_{0}=$ $-\frac{18b^{2}}{a_{\rm X}}\sum_{m>n}(\phi^{\alpha}_{m}\phi^{\alpha}_{n})^{2}$
$-\frac{36b^{2}}{a_{\rm X}}(c_{12}^{\alpha}c_{34}^{\alpha}+c_{13}^{\alpha}c_{24}^{\alpha}+c_{14}^{\alpha}c_{23}^{\alpha})\phi^{\alpha}_{1}\phi^{\alpha}_{2}\phi^{\alpha}_{3}\phi^{\alpha}_{4},$ (36)
and
$\delta F_{1}^{\alpha}=$ $-\frac{108b^{3}}{a_{\rm X}^{2}}\Bigg{\{}\sum_{m>n}\Big{[}u_{0}^{\alpha}\cos(2\theta^{\alpha}_{mn})+v_{0}^{\alpha}\sin(2\theta^{\alpha}_{mn})\Big{]}(\phi^{\alpha}_{m}\phi^{\alpha}_{n})^{2}$
$+6\Big{(}u_{0}^{\alpha}\cos\bar{\theta}^{\alpha}+v_{0}^{\alpha}\sin\bar{\theta}^{\alpha}\Big{)}\phi^{\alpha}_{1}\phi^{\alpha}_{2}\phi^{\alpha}_{3}\phi^{\alpha}_{4}\Bigg{\}},$ (37)
where $\theta_{mn}\equiv\theta_{m}+\theta_{n}$ and $\bar{\theta}^{\alpha}\equiv\theta_{1}^{\alpha}+\theta_{2}^{\alpha}+\theta_{3}^{\alpha}+\theta_{4}^{\alpha}$. Equation (36) renormalizes $c^{\prime}$ and $c^{\prime\prime\prime}$ as
$c^{\prime}\to c^{\prime}-\frac{72b^{2}}{a_{\rm X}},\quad c^{\prime\prime\prime}\to c^{\prime\prime\prime}-\frac{144b^{2}}{a_{\rm X}}.$ (38)
The terms appearing in $\delta F_{1}^{\alpha}$ are emergent fifth-order coupling constants between the mode at the $\Gamma$ and the L points. Since we have kept terms up to the fourth orders in the original free energy in Eq. (6), cases for large $108b^{3}/a_{\rm X}$ are beyond the applicable range of the approximation.
Appendix C Free energy for the orthogonal double-q configuration
[…]
$F=$ $\sum_{n=1,2}\Big{[}(a_{\rm L}+J_{\rm L}+\lambda_{c}^{(1)}\phi_{0}^{2})\phi_{n}^{2}+\frac{c}{2}\phi_{n}^{4}\Big{]}+\frac{c^{\prime}}{2}\phi_{1}^{2}\phi_{2}^{2}$
$-6\lambda_{b}\phi_{0}(\phi_{1}^{2}-\phi_{2}^{2})\cos(2\theta+\theta_{0})$
$+2\lambda_{c}^{(2)}\phi_{0}^{2}\Big{[}\phi_{1}^{2}\cos^{2}(\theta-\theta_{0})+\phi_{2}^{2}\sin^{2}(\theta-\theta_{0})\Big{]}$
$+\frac{216b^{3}}{a_{\rm X}^{2}}\phi_{0}\phi^{2}_{1}\phi^{2}_{2}\cos(4\theta-\theta_{0})+F_{\Gamma},$ (39)
where $c^{\prime}=2c-72b^{2}/a_{\rm X},\ \lambda_{b}=b,\ \lambda_{c}^{(1)}=\lambda_{c}^{(2)}=c$, $F_{\Gamma}$ is given by Eq. (10). Let us briefly discuss possible stationary angle $\theta$ under magnetic fields. For simplicity, we assume $\theta_{0}=0$, since $\tilde{h}$ induces $u_{0}$ under both $\bm{H}\parallel[001]$ and $[110]$. We treat $\phi_{0,1,2}$ as given parameters and the stationary condition of $F$ with respect to $\cos(2\theta)$ leads to
$\cos(2\theta)=\frac{6\lambda_{b}-\lambda_{c}^{(2)}\phi_{0}}{216b^{3}/a_{\rm X}^{3}}\left(\frac{\phi_{1}^{2}-\phi_{2}^{2}}{\phi_{1}^{2}\phi_{2}^{2}}\right).$ (40)
This means that $\phi_{1}/\phi_{2}\to 1$ for $\theta\to\pi/4$ unless $\phi_{0}$ is tuned to $6\lambda_{b}/\lambda_{c}^{(2)}$, which is the phase III2q seen in the main text. The free energy for the phase III2q is
$F_{{\rm III}_{2q}}=$ $2\big{[}a_{\rm L}+J_{\rm L}+(\lambda_{c}^{(1)}+\lambda_{c}^{(2)})\phi_{0}^{2}\big{]}\phi_{1}^{2}$
$+\left(c+\frac{c^{\prime}}{2}-\frac{216b^{3}}{a_{\rm X}^{2}}\phi_{0}\cos\theta_{0}\right)\phi_{1}^{4}+F_{\Gamma}.$ (41)
One can also see that the deviation $\delta\phi=\phi_{1}-\phi_{2}\neq 0$ from the phase III2q accompanies changes in the angle as $\theta=\pi/4+\epsilon$. The quadratic form of $\delta\phi$ and $\epsilon$ describe the stability of the phase III2q and when one of the eigenvalues vanishes a second-order transition takes place toward the phase II2q.
For $\theta=0$, i.e., the phase I2q, the free energy is given by
$F_{\rm{I}_{2q}}=$ $\sum_{n=1,2}\Big{[}(a_{\rm L}+J_{\rm L}+\lambda_{c}^{(1)}\phi_{0}^{2})\phi_{n}^{2}+\frac{c}{2}\phi_{n}^{4}\Big{]}+\frac{c^{\prime}}{2}\phi_{1}^{2}\phi_{2}^{2}$
$-6\lambda_{b}\phi_{0}(\phi_{1}^{2}-\phi_{2}^{2})\cos\theta_{0}$
$+2\lambda_{c}^{(2)}\phi_{0}^{2}\Big{[}\phi_{1}^{2}\cos^{2}\theta_{0}+\phi_{2}^{2}\sin^{2}\theta_{0}\Big{]}$
$+\frac{216b^{3}}{a_{\rm X}^{2}}\phi_{0}\phi^{2}_{1}\phi^{2}_{2}\cos\theta_{0}+F_{\Gamma}.$ (42)
Rotating $\theta$ leads to the transition to the phase II2q. Assuming $|\theta|\ll 1$ and $\theta_{0}=0$ in Eq. (39), we obtain the quadratic part of $F$ in $\theta$ as
$F=$ $\Big{[}12\lambda_{b}(\phi_{1}^{2}-\phi_{2}^{2})-\lambda_{c}^{(2)}\phi_{0}(\phi_{1}^{2}-2\phi_{2}^{2})-\frac{1728b^{3}}{a_{\rm X}^{2}}\phi^{2}_{1}\phi^{2}_{2}\Big{]}\phi_{0}\theta^{2}.$ (43)
Thus, a second-order transition to the phase II2q takes place when []delimited-[][$\cdots$][ ⋯ ] in Eq. (43) vanishes.